Find the complete integral of p+q=x+y


1
Expert's answer
2021-09-07T00:51:03-0400

Solution

Let’s consider the transformation  α = x – y,  β = y

Inverse transformation is x = α + β,  y = β

Let u(x, y) = v(α, β). Then, we have

p + q = ux + uy =  vα (∂α/∂x) + vβ (∂β /∂x) + vα (∂α/∂y)  + vβ (∂β /∂y)  =  vα - vα + vβ = vβ 

 

Therefore, the partial differential equation becomes

vβ =  α + 2β

Integrating this equation

v = α β + β2 + g(α) 

Here g(α) is an arbitrary function of α.

So u(x, y) = α β + β2 + g(α) = (x – y)y+y2 + g(x – y) = x y + g(x – y)

Answer

u(x, y) = x y + g(x – y), where g(α) is an arbitrary function of α.